OpenStudy (anonymous):

A 35.0 kg dolphin decelerates from 12.0 to 7.50 m/s in 2.30s to join another dolphin in play. What average force was exerted to slow him if he was moving horizontally? (The gravitational force is balanced by the buoyant force of the water)

OpenStudy (mrnood):

first answer these questions: 1) What is the total change in velocity 2) What is the time it takes to change (Ans= 2.3s) 3) What is the average change per second The answer to 3 gives you the average acceleration NOW: Force = m x a and you know the mass and you know the acceleration so you can calculate the force

OpenStudy (anonymous):

Force =(mass)(acceleration) F = (35.0kg)(2.3s) F = 80.5 is that right?

OpenStudy (mrnood):

no answer the 3 questions first - you need to calculate the acceleration from those questions....

OpenStudy (mrnood):

you wrote Force =(mass)(acceleration) - this is correct F = (35.0kg)(2.3s)

OpenStudy (mrnood):

Force =(mass)(acceleration) F = (35.0kg)(2.3s) - THIS is (mass)(time) you need to work out acceleration first

OpenStudy (mrnood):

acceleration = rate of change of velocity in this case average acceleration = change of velocity /2.3s

OpenStudy (anonymous):

a = V-Vo/ t a = 7.50m/s - 12.0m/s / 2.30s a = -1.957m/s^2 is that right?

OpenStudy (mrnood):

yes Now you can use F=ma to work out the average F

OpenStudy (anonymous):

f = (35.0kg)(-1.957) f = -68.495 is it right?

OpenStudy (mrnood):

Yes - but you should always include units if Mass is in kg and a is in ms^-2 then what is the unit of Force?

OpenStudy (anonymous):

-68.495Newton's

OpenStudy (mrnood):

OK - good

OpenStudy (anonymous):

Thanks MrNood. I really appreciate your help.

OpenStudy (mrnood):

np